OpenStudy (anonymous):

5/9w+8/9=3w (show work)

OpenStudy (anonymous):

\[5w/9 + 8/9 = 3w\]\[27w/9 - 5/9 = 8/9\]\[w = 4/11\]

OpenStudy (anonymous):

Thnx :)